Are the following trinomials perfect squares?
1. x^2-2x+1
2. x^2+4x-4
Find the constant term that should be added to make the following expression a perfect square trinomial.
x^2+8x
Find the constant term that should be added to make the following expression a perfect square trinomial.
x^2+3x

Answer by stanbon(75887)

Are the following trinomials perfect squares?
1. x^2-2x+1 = (x-1)^2; Yes
-----------
2. x^2+4x-4 ; no
-------------
Find the constant term that should be added to make the following expression a perfect square trinomial.
x^2+8x+(8/2)^2 or x^2+8x+16
---------------
Find the constant term that should be added to make the following expression a perfect square trinomial.
x^2+3x+(3/2)^2 = x^2+3x+(9/4)
----------------
Cheers,
Stan H.